OpenStudy (yacoub1993):

Multiply or divide. 5x+10/x+2 * 2x/4x-10

OpenStudy (yacoub1993):

\[\frac{ 5x+10 }{ x+2 }\times \frac{ 2x }{ 4x-10 }\]

OpenStudy (anonymous):

take 5 common from 5x+10 and 2 common from 4x-10 then divide and multiply..........

OpenStudy (yacoub1993):

still didnt get @RaviKiran can u slove further more

OpenStudy (anonymous):

it becomes( 5(x+2)/x+2)*(2x/2(2x-5))

OpenStudy (anonymous):



OpenStudy (yacoub1993):

owk then

OpenStudy (yacoub1993):

is that it ?

OpenStudy (anonymous):

ha thats all aftr cancillation of terms multipy the remaining and u vl get it

OpenStudy (yacoub1993):

thnxs @RaviKiran